OpenStudy (noahbred):

Please confirm its -8 Solve. z + 17 = 9 A. 9 B. 8 C. –9 D. –8

jimthompson5910 (jim_thompson5910):

Let's check z + 17 = 9 -8 + 17 = 9 ... replace every `z` with `-8` 9 = 9 ... true Since the last equation is true, the first equation is true when z = -8. So the solution is confirmed.

OpenStudy (noahbred):

Oh thankss... :D

jimthompson5910 (jim_thompson5910):

no problem